OpenStudy (anonymous):

Use the given formula to solve the following problem: When the amount financed is $15,480, the monthly payment is $621.33, and the monthly interest rate is 0.00583, how long do you have to pay? A. 18 months B. 23 months C. 27 months D. 32 months

OpenStudy (anonymous):

log[P/(P-Ar)] n=________________ log(1+r)

OpenStudy (anonymous):

@petiteme

OpenStudy (anonymous):

petiteme do u know how to do this???

OpenStudy (wolf1728):

I could solve this (I think) In that formula, what does 'Ar' mean?

OpenStudy (anonymous):

A IS amount and r is rate

OpenStudy (wolf1728):

I'm trying to use a formula I ise in my own online calculator: time = log(1+(rte/(pmt/(amt)-rte))) / Math.log(1+rte))/12; And what does amount rate mean?

OpenStudy (anonymous):

amount is the price of the car ,house or idk but is what really cost and the rate is the interests the we have to pay like taxes

OpenStudy (wolf1728):

Well, I will have to leave for a few hours, but I'll come back to this.

OpenStudy (anonymous):

okay thanks

OpenStudy (wolf1728):

Okay, I have the answer and I used my formula. Principal = $15,480 monthly payment = $621.33 monthly interest rate = 0.00583 months = log(1+(rate/(pmt/(princ)-rate))) / log(1+rate) months = log(1 + (.00583 / (621.33 / (15,480) - .00583))) / log ( 1.00583) months = log(1.169933208) / log(1.00583) months = 0.0681610684 / 0.0025245848 months = 26.9989221736 I'd guess 27 is the answer considering that does round neatly to 27 and that choice C = 27. There it is. I have no idea how to use that other formula but at least you have an answer.